A basketball player scored 27

times during one game. He

scored a total of 49

​points, two for each​ two-point shot and one for each free throw. How many​ two-point shots did he

​make? How many free​ throws?

Answer by Boreal(15235)

x=2 point shots
y=free throws
x+y=27
2x+y=49
y=-2x+49
substitute into first
x-2x+49=27
-x=-22
x=22 two point shots (44 points)
y=5 free throws (5 shots)
27 times scoring for 49 points.